OpenStudy (anonymous):

Please help. Two boys want to use a seesaw, and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (blackbird02):

You just multiply both sides of the equation by 1/w2 to cancel out w2 so that d2 would be the only variable on one side of the equation \[(\frac{ 1 }{ w2})(w1 • d1)=(w2 • d2)(\frac{ 1 }{ w2})\] After you cancelled out w2, this would be the resulting equation, \[d2=\frac{ w1 •d1 }{w2 }\]